In a study of the formation of NOx air pollution, a chamber heated to 2200°C was filled with air (0.790 atm N2, 0.210 atm O2). What are the equilibrium partial pressures of N2, O2, and NO if Kp = 0.0460 for the following reaction:

N2+O2 (equilibrium arrow) N2O2

Part 1

P(N2) = atm

Part 2

P(O2) = atm

Part 3

P(NO) = atm
Answer #1

N2 + O2 -----------> 2NO , Kp = 0.046

0.79 0.21 0 (at t=0)

0.79-x 0.21-x 2x (at t=equilibrium)

Kp = (2x)^2 / (0.79-x)*(0.21-x)

0.046 (0.79-x) (0.21-x) = 4x^2

0.046 (x^2 - x + 0.1659) = 4x^2

0.046x^2 - 0.046x + 0.0076314 = 4x^2

3.954x^2 + 0.046x - 0.0076314 = 0

After solving above equation, we get

x = 0.0385

So, P(NO) = 2x = 2*0.0385 = 0.077 atm

P(N2) = 0.79-x = 0.79 - 0.0385 = 0.7515 atm

P(O2) = 0.21 - x = 0.21 - 0.0385 = 0.1715 atm ..

. ANSWER